# Projective Modules and Cohomology for Integral Basic Algebras

### v1

@@ 13. Connected CW Left Regular Bands satisfy D
In this section, we consider the class of algebras $kB$ where $B$ is a left regular band, i.e., $B$ is a semigroup satisfying the identities $x^{2}=x$ and $xyx=xy$ for all $x,y\in B$. Since $B$ may not have an identity, the algebra $kB$ does not necessarily admit a unit. However, following Margolis–Saliola–Steinberg [24], to guarantee a unit, it suffices to assume that $B$ is a connected CW left regular band – we shall not need the full definition here and refer the reader to their paper. In the case $B$ is a monoid, it is also $\mathcal{R}$-trivial (see Section 12). This class of algebras includes the face algebras of hyperplane arrangements in real spaces. In the special case when the hyperplane arrangement is given by the reflection arrangement of a (finite) Coxeter group $W$, the descent algebra of $W$ is isomorphic to opposite of the fixed point space of the face algebra under the action of $W$ as shown by Bidigare [8].
Throughout this section, we assume that $B$ is a connected CW left regular band, $\Lambda(B)$ the support semilattice of $B$ and
@@
the support map which is a semigroup homomorphism. Recall that, given a graded (finite) poset $(\leq,P)$, the closed interval $[X,Z]$ is the subset $\{Y:X\leq Y\leq Z\}$ of $P$. Similarly, we denote $(X,Z)$ for the open interval. The rank (or length) $\ell([X,Z])$ of $[X,Z]$ is the length of a longest chain starting at $X$ and ending at $Y$.
Theorem 13.1 ([24, Theorem 1.1(4,5)]).
@@
Consider the face algebra ${\mathcal{O}}\mathcal{F}$ with the hyperplane arrangement defined by the hyperplanes $H_{1},\ldots,H_{n}$ in $\mathbb{R}^{d}$. Suppose further that there exists a subspace $X$ such that $H_{i}\cap H_{j}=X$ for all $1\leqslant i\neq j\leqslant n$. By [28, Proposition 8.5], the Ext quiver $Q$ of ${\mathbb{F}}\mathcal{F}$ (or $\mathbb{K}\mathcal{F}$) is given by the diagram below.
@@
In general, does ${{\mathcal{O}}B}$ satisfy E?

### v2

@@ 13. Connected CW Left Regular Bands satisfy D
In this section, we consider the class of algebras $kB$ where $B$ is a left regular band, i.e., $B$ is a semigroup satisfying the identities $x^{2}=x$ and $xyx=xy$ for all $x,y\in B$. Since $B$ may not have an identity, the algebra $kB$ does not necessarily admit a unit. However, following Margolis–Saliola–Steinberg [25], to guarantee a unit, it suffices to assume that $B$ is a connected CW left regular band – we shall not need the full definition here and refer the reader to their paper. In the case $B$ is a monoid, it is also $\mathcal{R}$-trivial (see Section 12). This class of algebras includes the face algebras of hyperplane arrangements in real spaces. In the special case when the hyperplane arrangement is given by the reflection arrangement of a (finite) Coxeter group $W$, the descent algebra of $W$ is isomorphic to the opposite of the fixed point space of the face algebra under the action of $W$ as shown by Bidigare [9].
Throughout this section, we assume that $B$ is a connected CW left regular band. Let $\Lambda(B)=\{Bb:b\in B\}$ be the support semilattice of $B$ and
@@
be the support map, defined as $\sigma(a)=Ba$, which is a semigroup homomorphism. Recall that, given a graded (finite) poset $(\leq,P)$, the closed interval $[X,Z]$ is the subset $\{Y:X\leq Y\leq Z\}$ of $P$. Similarly, we denote $(X,Z)$ for the open interval. The rank (or length) $\ell([X,Z])$ of $[X,Z]$ is the length of a longest chain starting at $X$ and ending at $Y$.
Theorem 13.1 ([25, Theorem 1.1(4,5)]).
@@
Consider the face algebra ${\mathcal{O}}\mathcal{F}$ with the hyperplane arrangement defined by the hyperplanes $H_{1},\ldots,H_{n}$ in $\mathbb{R}^{d}$. Suppose further that there exists a subspace $X$ such that $H_{i}\cap H_{j}=X$ for all $1\leqslant i\neq j\leqslant n$. By [30, Proposition 8.5], the Ext quiver $Q$ of ${\mathbb{F}}\mathcal{F}$ (or $\mathbb{K}\mathcal{F}$) is given by the diagram below.
@@
Let $B$ be a connected CW left regular band. Does ${{\mathcal{O}}B}$ satisfy E?